OpenStudy (anonymous):

How to find the derivative of 1 - (1/x) using power rule

OpenStudy (cwrw238):

this is 0 - -1x^-2 = 1/x^2

OpenStudy (cwrw238):

i'm assuming the power rule is d(ax^n) dx = anx^(n-1)

OpenStudy (anonymous):

oh ok, thanks

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

derivative of \[y=\frac{1}{x}\] is \[y'=-\frac{1}{x^2}\]easier just to remember than use the power rule

OpenStudy (anonymous):

yea i thought u just leave the 1 as 1 cuz for like 3x^2 u leave the 3 as 3... so i got confused about it :(

OpenStudy (anonymous):

ah that is the difference between a coefficient and a constant. or the difference between additon and multiplication. think of it this way, y = 1 is a line with slope 0, so the constant contributes nothing to the derivative

OpenStudy (anonymous):

so when there is an addition or subtraction, say 3 - 2x^2 the 3 is not a constant anymore , and have to take the derivative of that right? so it will be 0 - 4x = -4x and the -2 is the constant

OpenStudy (anonymous):

when you take the derivative of the number with no variable attached, you get zero

OpenStudy (anonymous):

so yes, you would get \[-4x\]

OpenStudy (anonymous):

oh ok this is clear thank you satellite :)